OpenStudy (anonymous):

how do you go about solving 2 equations with 3 unkowns!?!? please kindly help . -3a + 4b + c = 0 6a - 2b + c = 0

OpenStudy (cwtan):

1st equation minus 2nd equation it will be 2 unknown left

OpenStudy (ash2326):

Assume any one of a or b or c and then find the other two

OpenStudy (anonymous):

rearrange the equations

OpenStudy (precal):

aren't we missing a third equation?

OpenStudy (anonymous):

c=3a-4b c=2b-6a

OpenStudy (anonymous):

ohhh ! i gotit already ! thank you people ~!!

OpenStudy (anonymous):

ehh @studymathlol then witht that how find ??

OpenStudy (anonymous):

lol since c = 3a-4b and c also =2b-6a there fore c = c and 3a-4b = 2b+ 6a

OpenStudy (anonymous):

geddit?

OpenStudy (anonymous):

yeap and then??

OpenStudy (cwtan):

and then find a in term of b sub into both equation

OpenStudy (anonymous):

how to get either a or b next?

OpenStudy (anonymous):

will come back to same eq.,..

OpenStudy (anonymous):

switch here switch there and you get 3a+6a=2b+4b 9a=6b

OpenStudy (anonymous):

yes and i tried finding a or b but subbing. i couldnt....=[

OpenStudy (anonymous):

precal is right we are missing a 3rd equation are you sure that is the full question? :/

OpenStudy (anonymous):

I think we need the third equation

OpenStudy (anonymous):

im not sure but i'm doing vectors chapters of alevel maths paper 3 questions.~

OpenStudy (anonymous):

trynna find equation from those vectors

OpenStudy (anonymous):

3 positiion vectors, find eq connecting all three

OpenStudy (anonymous):

yah or else you will get